OpenStudy (anonymous):

Area with sums problem: Find the upper and lower sums of f(x) = -2x + 4 on the interval [0, 2] using four subintervals of equal length. 'Note: do not use the summation formulas; calculate the area of the rectangles directly.'

OpenStudy (anonymous):

ok so if i can't use summation formulas, how should i go about this? i'll get to work drawing a graph

OpenStudy (anonymous):

 something like this? width for each should be 0.5 (b-a/n= 2-0/4= 0.5)